OpenStudy (anonymous):

solve the equation sr(x-2) = 4

OpenStudy (anonymous):

x-2=4^2 x-2=16 x=18

OpenStudy (anonymous):

but dont you sr 18?

OpenStudy (anonymous):

what's "sr" in the equation?

OpenStudy (anonymous):

no we dont need to.. because we want to find out x... not sr x..

OpenStudy (anonymous):

sr means sqaureroot

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

then @vedic is correct.

OpenStudy (anonymous):

then how would you do this one sr(x) +1=5

OpenStudy (anonymous):

sr x=5-1 srx=4 x=4^2 x=16

OpenStudy (anonymous):

and you dont have to square root that?

OpenStudy (anonymous):

no we dont have to...